OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. -6x + 3y = 18 4x - 3y = 6 2x + 4y = 24 6x + 12y = 36 3x - y = 14 -9x + 3y = -42 5x + 2y = 13 -x + 4y = -6

OpenStudy (precal):

look at your pair of equations, look for a pattern for example 2x + 4y = 24 6x + 12y = 36 2x times 3 is 6x 4y times 3 is 12y 24 times 3 is what? if it is 36 then this is your solution, if not then it is not

OpenStudy (anonymous):

72 so it's not

OpenStudy (anonymous):

Im going to go with c? 3x-y=14 -9x+3y=-42 3×-3=-9 -1×-3=3 14×-3=-42

OpenStudy (precal):

yes

OpenStudy (anonymous):

okay thanks :)

OpenStudy (precal):

np